OpenStudy (anonymous):

Please help me find a common denominator for 1/sqrt(9+h) and 1/sqrt9

OpenStudy (zepp):

It should be sqrt(81 + 9h)

OpenStudy (zepp):

Multiply sqrt(9+h) and sqrt (9)

OpenStudy (anonymous):

That isnt right..

OpenStudy (anonymous):

let me get someone to help ya quick

OpenStudy (anonymous):

thanks for coming hero

hero (hero):

I just hate when it takes so damn long for the page to load.

OpenStudy (anonymous):

I may have asked the wrong question. I have to subtract those two fRACTIONS from one another I may be asking for help with the wrong thing. The equation is 1/sqrt(9+h)-1/sqrt9 all over h and I am solving for h. How do I get rid of the square roots so I can find what h is?

hero (hero):

Well, sqrt{9} = 3 for starters

OpenStudy (anonymous):

sqrt9-h over 9-9+h?

hero (hero):

@mertsj you should respectfully allow me to continue helping since I have already begun. I know you're probably going to post a full solution, but I was in the middle of helping the person with it step by step. You should do what people like Fool for Math do and just continue to allow me to help the person. There are plenty other questions you could be helping others with.

OpenStudy (mertsj):

Well, except that the asked asked me for help.

hero (hero):

Why did you delete your post then?

hero (hero):

You could have just said that rather than delete it

OpenStudy (anonymous):

how is the denominator sqrt (9-h)*(9+h)?

hero (hero):

It's called conjugation

OpenStudy (mertsj):

I would like to help you but I must wait until Hero is done.

hero (hero):

Go ahead Mertsj. You help people by posting the full solutions rather than guiding them throught the problem. If that's how you want to help, then by all means go ahead.

OpenStudy (anonymous):

can someone help me figure out what to do after i have that fraction?

hero (hero):

I also would like to help, but I have to deal with @mertsj interrupting my explanations.

OpenStudy (anonymous):

he is gone now.

OpenStudy (anonymous):

please help me?

hero (hero):

Actually, just multiply the first fraction by \[\frac{\sqrt{9+h}}{\sqrt{9+h}}\]

OpenStudy (anonymous):

okay I have sqrt(9+h)over 9+h - 1/3 all over h?

hero (hero):

There's no minus. We're adding the fractions. Sorry if I confused you earlier

hero (hero):

Now I may have no choice but to post the full solution.

OpenStudy (mertsj):

\[\frac{1}{\sqrt{9+h}}+\frac{1}{\sqrt{9}}=\frac{1}{\sqrt{9+h}}+\frac{1}{3}\]=

OpenStudy (anonymous):

original equation is minus 1/sqrt 9

hero (hero):

For some reason everyone else thought it was plus

OpenStudy (mertsj):

\[\frac{3}{3\sqrt{9+h}}-\frac{\sqrt{9+h}}{3\sqrt{9+h}}=\frac{3-\sqrt{9+h}}{3\sqrt{9+h}}\]

OpenStudy (anonymous):

now I am just confused :(

hero (hero):

There's still a square root in the denominator, but okay @mertsj

OpenStudy (mertsj):

I do believe that the problem asked to find the common denominator for the two fractions. as I have shown, the common denominator is 3sqrt(9+h) So if you disagree or are confused, please clarify exactly what the problem is. To rationalize the denominator or to find the common denominator. In any case I have other things to do that fuss about such stupidity.

hero (hero):

Well, I see your point, @mertsj, however, if you rationalize, you end up with a common denominator of 3(9+h) or 27 + 3h

OpenStudy (anonymous):

\[f \prime(a)= \lim_{h \rightarrow 0} (1/\sqrt{9+h})-1/\sqrt{9})/h\] is where I need to end up I need the steps to easily sub in 0 for h so that I do not have a 0 in the denominator.

hero (hero):

You should have just posted that to begin with

OpenStudy (anonymous):

i did in a previous problem and no one helped me so I thought I would try to get help with the steps I knew came next in smaller easier problems. But IF you could help me with this as is it would be greatly appreciated :)

hero (hero):

Then afterwards, you could have simply stated, "I'm having trouble finding a common denominator".

hero (hero):

I will try to help

OpenStudy (anonymous):

i just basically need to get the h out of the denominator

hero (hero):

Yes, it's actually simpler than you think. I will write it out

OpenStudy (anonymous):

okay thank you

hero (hero):

Well, it wasn't as simple as I originally thought but it is doable. It essentially comes down to remembering distribution and simplification rules

hero (hero):

I see why you're having difficulty with this. Again I want to stress posting the complete problem the first time. It would have saved everyone a lot of time.

hero (hero):

Just to confirm: \[f'(a) = \lim_{h \rightarrow 0} \frac{\frac{1}{\sqrt{9+h}} - \frac{1}{\sqrt{9}}}{h}\] Is that correct?

OpenStudy (anonymous):

you guys are still at it?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

congrats on the thoroughness hero!

hero (hero):

Only because at the last minute she posts the actual problem she's working with.

OpenStudy (anonymous):

oh, lol

OpenStudy (anonymous):

im sorry.. jeez.

hero (hero):

Not to worry. I will get you the solution

hero (hero):

You will find the solution there. Sorry for the trouble.

OpenStudy (anonymous):

the all-mighty wolfram!

hero (hero):

I wonder if you had to find the solution the old fashioned way

hero (hero):

Because you wrote f'(a)

hero (hero):

A proper solution will be posted before I go to bed.

OpenStudy (anonymous):

wait so what you posted isnt right?

hero (hero):

f'(a) = -1/54 is correct It's just that the solution steps that wolfram alpha posted aren't the proper steps for posting f'(a) using the definition of a derivative which is what you are working on.

OpenStudy (anonymous):

oh.. well how do I do that?

hero (hero):

Well, I checked with a personal source I have (calculus expert) and even they admit that it is difficult to find an approach that would cancel the fraction 1/h

OpenStudy (anonymous):

:'(

OpenStudy (anonymous):

okay so i really need help i have class at 6 which is in one hour @Mertsj can u help?

hero (hero):

I finally have the solution and I will be posting it shortly.

OpenStudy (anonymous):

okay thank you very much!!

OpenStudy (anonymous):

thank you sooo much you saved my butt!!

OpenStudy (anonymous):

where did u go to get that help @Hero

OpenStudy (zepp):

Wow that was brillant!

hero (hero):

What makes you think I went to anybody? The only thing that I did differently was instead of trying to conjugate the denominator, I decided to conjugate the numerator. I would have done the steps on here, but it lags too much and plus using that drawing board is just awful.

hero (hero):

On top of that, you should have just posted the original problem from the beginning and it would have been solved in 10 minutes at most.

OpenStudy (anonymous):

it was just a question sorry

OpenStudy (anonymous):

how did you get the -h in the 4th step from the bottom?

OpenStudy (anonymous):

nvmfigured it out